OpenStudy (anonymous):

a gardener plants 200 flowering plants of the same species. 98 plants have white (v) flowers and the remaining have violet (v) flowers. what must be the allelic frequency of each type of flower? (p+p=1, p2+ 2pq+ q2=1)

OpenStudy (anonymous):

A VIOLET FLOWERS = 0.55 WHITE FLOWERS = 0.45 B VIOLET FLOWERS = 0.62 WHITE FLOWERS = 0.38 C VIOLET FLOWERS = 0.3 WHITE FLOWERS = 0.7 I THINK ITS C

OpenStudy (carolina4567):

The first thing to do is figure out the frequency of the white and violet flowers. You have 200 total, so: white = number white / number total violet = (number total - number white) / number total Get your data into a form where you can start using the Castle Hardy Wienburg relations on it...

OpenStudy (anonymous):

AM STILL KINDA CONFUSED ABOUT IT ;(

OpenStudy (carolina4567):

Look theres 200 flowers 98 are white 200 -98 = ?

OpenStudy (anonymous):

ahhhhhhh thank you

OpenStudy (carolina4567):

102 right?

OpenStudy (carolina4567):

sooooo we have 98 white and 102 violet correct ?

OpenStudy (anonymous):

yes

OpenStudy (carolina4567):

98/200 =

OpenStudy (carolina4567):

0.49

OpenStudy (carolina4567):

200 - 98 = 102 / 200 = 0.51

OpenStudy (carolina4567):

So most logical answer is ?

OpenStudy (anonymous):

yeah i got it right thanks alot

OpenStudy (carolina4567):

Np anytime

OpenStudy (carolina4567):

Become a fan so u know when im on